Required Information [The following information applies to the questions displayed below.] Part 1 of 3 In 2018, Laureen is cub. Laureens AGI IS $95.000. American opportunity credit Lifetime learning creditc. Laureens AGI IS $45,000 and Laureen paid $12,160 (not $1,780) for Ryan to attend graduate school (l.e, his fifth year, no
Answer #1
Answer:-
a) Laureen's AGI - $45,000
AOTC is AMERICAN OPPORTUNITY TAX CREDIT
For 2 daughter - AOTC is - (2000*2child)+(800*25%+2child)
=4,000+400
=4,400
For Ryan - 1780
AOTC =4400+1780
AOTC =6,180
Laureen lifetime learning credit - Eligible is 2000 (The amount of the credit is 20 percent of the first $10,000 of qualified education expenses or a maximum of $2,000 per return)
so in above case it is - 1280*20% =256 (Since AGI is below clip of 56000 he can claim same)
So, 6,180+256 = 6,436 is eligible deduction
b)
Since AGI is 95000
AOTC can't be calimed if AGI is above 90000 and hence AOTC is zero and Lifetime learning credit can't be claimed if AGI is above 56000.. Hence it is zero education credit
c)
For Daughter it is same as a above i.e. 4,400
For Ryan it is = 2000+(10000*25%) or maximum 4000
=2000+2500 or 4000
so 4000 is allowed
so AOTC total of 8400 and LLC of 256 so claimed is 8,656